OpenStudy (anonymous):

integrate cos^3t sin^4t

OpenStudy (anonymous):

use u-substitution

OpenStudy (anonymous):

@RolyPoly

OpenStudy (anonymous):

\[\int cos^3t sin^4tdt\]\[=\int cost(cos^2t) sin^4tdt\]\[=\int(cos^2t) sin^4td(sint)\]\[=\int(1-sin^2t) sin^4td(sint)\]\[=\int (sin^4t-sin^6t)d(sint)\]\[=...\]

OpenStudy (anonymous):

den it become..sin^5t - sin^7t

OpenStudy (anonymous):

how do we solve that ? :P...sorry i am totaly weak in this :( @RolyPoly

OpenStudy (anonymous):

My bad! Use u-sub then... Let u=sint What is du/dt?

OpenStudy (anonymous):

cost

OpenStudy (anonymous):

Yup. So, du = cost dt cos^3t = cost (cos^2 t) Use identity sin^2 t + cos^2 t = 1 cos^2 t can be rewritten as 1-sin^2 t Understand so far?

OpenStudy (anonymous):

i guess you made a mistake at the previous derivation...please check !

OpenStudy (anonymous):

May I know where I made mistakes?

OpenStudy (anonymous):

∫cost(cos^2t)sin4tdt ∫(cos^2t)sin4t(sint) where did the cost t go ?

OpenStudy (anonymous):

The dt changed to d(sint) . Since d(sint) = cost dt

OpenStudy (anonymous):

oh!! ...

OpenStudy (anonymous):

Got it?

OpenStudy (anonymous):

it then becomes integrate of U^5t-U^7t

OpenStudy (anonymous):

Not really. ∫(cos^2t)sin^4t d(sint) =∫(1-sin^2t)sin^4t d(sint) =∫sin^4t - sin^6t d(sint) Let u = sint =∫u^4 - u^6 du =... ?

OpenStudy (anonymous):

and then..u^5/5 - u^7/7

OpenStudy (anonymous):

Yes. And put back u=sint into that one. u^5/5 - u^7/7 = ... ?

OpenStudy (anonymous):

awesomenss...:D :D ..and i have another one..integration is so hard :(

OpenStudy (anonymous):

Couldn't agree more!